NAD 2200 current draw rises

So the laser thermometer is great ( a new toy).


Initially I could not see any difference between the channels, so I reverted to replacing the active components:
Q326/Q328- no difference
Small signal transistors- all the C2240 and A970 before the bias section- no difference
then the bias transistor Q334- now the bias was way high, dropped this back and back and the rise in AC draw slowed. Then checked the bias voltages (again) and found the +ve side a bit high at 1.89v- so replaced D310 and D302/D304- so now it is 1.8v (should be 1.77v) and the -ve side is -1.75v.


so now the amp starts cold at AC draw 0.4A (correct) and slowly increases to 0.5A. Then used the temp laser and R338 on the Left (good) channel is about 40c whereas on the right it is 20c- which suggests that Q324 and/or Q328 and Q332 are not operating as they should, that is not pulling any current.


Q328 has been replaced as has Q324, ZD308 measures correctly- so I will focus on this section.


I just want to be sure that the underlying problem is solved
 
so to round out this thread:


I replaced R346, D328 and C318 and the AC consumption is hovering at 0.5A, which is OK.



I then replaced Q332 with another 2SC3281. The one I removed tested fine and had an HFE of 70, whereas these generally are around 50, as the one I replace it with has.



As I adjust the idle to 7.5mv, the AC current seems stable. I will run it for the next couple of days to make sure it behaves.


Just a note as well- all of the transistors and diodes (except the initial fried ones) tested good (peak tester), and yet when replaced the instability has effectively gone.


Thanks for all the help it is much appreciated.